OpenStudy (anonymous):

Evaluate the formula A = p2 p=22/7 and r=7/4

OpenStudy (anonymous):

A=p2 A=22/7(2) A=44/7

OpenStudy (anonymous):

Hello are you there. Here's how this program works. If you get help from someone, you give them credit by giving them medal. They will do they same for you if you have the best response on a question, so did my answer help?

OpenStudy (anonymous):

@momzoo?